OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 1^2 + 4^2 + 7^2 + . . . + (3n - 2)2 = (n(〖6n〗^2-3n-1))/2

OpenStudy (anonymous):

whoever solves will get medal

OpenStudy (anonymous):

lol idk im new to this thing, can someone plz just help me

OpenStudy (anonymous):

@kaos_gabz help please

OpenStudy (anonymous):

S1 = 1 S2 = 17 s3= 67

OpenStudy (anonymous):

can u show me how to do it plz, i wanna understand

OpenStudy (anonymous):

what he said..

OpenStudy (anonymous):

S1 is sum of 1 term = 1^2 =1 S2 is sum of 2 terms = 1^2 + 4^2 =17 you do S3

OpenStudy (anonymous):

s3=3 terms 1^2+4^2+7^2=67

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but how do i show that each statement is true

OpenStudy (anonymous):

like question is asking

OpenStudy (anonymous):

read my comment above, it has 2 parts : "statement" and "proof" "statement" is :" S1 is sum of 1 term, it is one" "proof" : S1= 1^2 =1 do the same with S2, S3

OpenStudy (anonymous):

i got it thx